Requirement 3 Either print a copy of your worksheet or make a copy of the worksheet in your workbook before proceeding You will need to refer back to this worksheet. Change the percentage completion with respect to conversion for the beginning inventory from 50% to 30%, but keep everything the same as in Requirement 2, The data area of your worksheet should now look like this 3 Data 4 Beginning work in process inventory: 5 Units in process 6 Completion with respect to materials 7 Completion with respect to conversion 8 Costs in the beginning work in process inventory: 9 10 Conversion cost 11 Units started into production during the period 12 Costs added to production during the period 13 Materials cost 14 Conversion cost 15 Ending work in process inventory: 16 Units in process 17 Completion with respect to materials 18 Completion with respect to conversion 800 85 % 30% S31,078 S34,840 17,600 Materials cost $798,842 $1,549,940 500 60% 80 % If your formulas are correct, you should get the correct answers to the following questions. (a) What is the equivalent units of production for materials? Equivalent units of production for materials
Answer #1

a) Equivalent unit of material = 17900+(500*60%) = 18200 unit

b) Equivalent unit of conversion = 17900+(500*80%) = 18300 Unit

c) Cost per equivalent unit of material = (31078+798842)/18200 = 45.60 per equivalent unit

d) Cost per equivalent unit of conversion = (34840+1549940)/18300 = 86.60 per equivalent unit

e) Cost of units transferred out = (45.60+86.60)*17900 = 2366380